How is three-fourths x minus 45=0 become 0?
The full equation is 11/12 x-12=1/6x+33

Answer by Alan3354(69443)

How is three-fourths x minus 45=0 become 0?
(3/4)x - 45 = 0
(3/4)x = 45
(1/4)x = 15
x = 60
-----------------
The full equation is 11/12 x-12=1/6x+33
------------
11x/12 - 12 = x/6 + 33
Multiply by 12 to eliminate fractions
11x - 144 = 2x + 396
9x = 540
x = 60